# CAMPOS II. The onset of protostellar disk substructures and planet formation

### v1

@@ 6 Conclusion
In summary, our ALMA study reveals that by $T_{\rm bol}\sim 200-400\,K$ protostellar disk substructure detection rates increased sharply to $\sim$60 %. Our statistical analysis confirms that substructures similar to those in Class II disks are already common by the Class I stage. We argue that this emergence is likely only an upper age limit, as our study suggests increasing optical depth toward the earliest stages.
The prevalence of substructures in the early protostellar stage suggests that planetesimals and giant planets form much more efficiently and earlier than predicted by traditional models (e.g., Pollack et al. (1996)). This early planet formation scenario is also supported by a low abundance of water in the inner Solar System (Morbidelli et al., 2016; Hyodo et al., 2019), the existence of several isotopic reservoirs in the Solar system (Kruijer et al., 2017) as well as the distribution of the calcium-aluminum-rich inclusions (Desch et al., 2018; Brasser & Mojzsis, 2020). Additionally, older protoplanetary disks (around Class II sources) might not have enough dust mass to form planetary systems (Manara et al., 2018).
Our study has pushed the onset of planet formation to the much earlier protostellar phase, in $\sim$0.2–0.4 Myr disks. During this time period, the central protostar still has a high accretion rate and is building up its mass.
@@
All data used in this study are public and can be downloaded from the ALMA Science Archive.
Software
Astropy: Greenfield et al. (2013); Astropy Collaboration et al. (2018, 2022), CASA: McMullin et al. (2007) SciPy: Virtanen et al. (2020)
@@ Acknowledgements.
C.H.H. and H.A are supported in part by NSF grants AST-1714710. C.H.H. is supported by NASA Hubble Fellowship Program under award HF2-51556. D.S.C. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2102405. Special thanks to e-HOPS group, Dr. Riwaj Pokhrel, and Dr. Thomas Megeath for providing the unpublished e-HOPs catalog for protostar classification. C.H.H. thanks Dr. Sean Andrews for the suggestion of considering the effective resolution in constructing a homogeneous sample for this statistical study. This paper makes use of the following ALMA data: ADS/JAO. ALMA #2016.1.00484.L, #2018.1.00028.S, #2019.1.00261.L, #2019.1.01792.S. ALMA is a partnership of the ESO (representing its member states),NSF(USA),and NINS (Japan), together with the NRC (Canada), NSC and ASIAA(Taiwan), and KASI(Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by the ESO, AUI/NRAO, and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.
@@ Appendix A Supplementary Materials
In Table LABEL:table:substructure_data, we present all the disks in this study complied from the CAMPOS survey (Hsieh et al. 2024), eDisk survey (Ohashi et al. 2023), DSHARP survey (Andrews et al. 2018a), and ODISEA survey (Cieza et al. 2021, 2019). All disks have a bolometric temperature less than 1900 K, and inclination angles less than $75^{\circ}$. The Table consists of a homogeneous survey of 5 molecular clouds, Corona Australis, Chamaeleon I & II, Ophiuchus & Ophiuchus North, with additional sources in Taurus, Lupus I and BHR 71 from the literature.
In Table LABEL:table:CAMPOS_Tbol, we cross-matched all the Corona Australis, Chamaeleon I & II, Ophiuchus North, and Ophiuchus sources in our CAMPOS survey with the Young Stellar Objects catalog from the Spitzer Space Telescope “cores to disks” (c2d) and “Gould Belt” (GB) Legacy surveys (Dunham et al. 2015), as well as the Extension of HOPS Out to 500 ParSecs (eHOPS) catalog to obtain the bolometric temperature ($T_{\rm bol}$), which serves as a proxy for relative evolutionary age for embedded protostellar systems. The eHOPS catalog contains 1-850 $\mu$m SEDs assembled from 2MASS, Spitzer, Herschel, WISE, and JCMT/SCUBA-2 data. It represents the latest and most reliable SED fitting to date. The first eHOPS paper, covering the Serpens and Aquila molecular clouds, was published by Pokhrel et al. (2023). For all other clouds, the SED and the protostellar system properties are available in NASA/IPAC Infrared Science Archive.
@@
CFHTWIR-Oph 43 Flat 440 570 0.12 0.38 4.4 144 2.0 NA 130 x 100 16.4 23.9 25.3 1
Oph-emb 23 Flat 440 570 0.12 8.68 50.8 144 46.2 NA 130 x 100 16.4 7.4 25.3 1
Oph-emb 26a Flat 500 620 0.97 1.9 27.8 144 10.1 NA 130 x 100 16.4 6.3 42.7 1
Oph-emb 26b Flat 500 620 0.97 1.89 27.7 144 10.1 NA 130 x 100 16.4 6.3 42.7 1
F-MM7 Flat 570 690 0.72 1.17 27.0 144 6.2 NA 130 x 100 16.4 5.2 39.6 1
Oph-emb 27 Flat 570 690 0.72 13.0 24.9 144 69.0 NA 130 x 100 16.4 25.2 39.6 1
Oph-emb 28 Flat 610 720 1.1 2.13 10.3 144 11.3 NA 130 x 100 16.4 24.3 44.0 1
@@
ISO-ChaI 101 Flat 650 820 0.09 12.06 73.8 179 41.5 C 99 x 65 14.4 6.0 23.6 1
V* DK Cha II 660 1500 30.0 210.59 56.7 181 708.7 NA 146 x 82 19.8 68.2 100.6 1
V* HO Cha a Flat 710 900 0.97 4.35 20.4 179 15.0 NA 120 x 90 18.6 14.8 42.7 1
V* HO Cha b Flat 710 900 0.97 4.43 22.0 179 15.3 NA 120 x 90 18.6 13.5 42.7 1
ISO-ChaI 204 II 720 1100 0.18 0.46 11.9 179 1.6 NA 120 x 90 18.6 7.2 28.0 1
ISO-ChaI 207 II 1100 1500 0.09 8.75 90.7 179 30.1 NA 100 x 65 14.4 4.4 23.6 1
ISO-ChaI 237 II 1100 1600 0.37 8.9 25.8 179 30.6 NA 120 x 90 18.6 17.7 33.5 1
Ass Cha T 1-15 II 1400 1700 0.45 12.7 46.5 179 43.7 NA 100 x 65 14.4 10.1 35.2 1
@@
101010 CFHTWIR-Oph 79 is also known as GY 263 or SKS 3-48 (Wilking et al. 2015; Narayanan et al. 2023). The classification of this source is debated. Narayanan et al. (2023) and Allen et al. (2002) classified this source as a Class II disk, while Wilking et al. (2015) classified it as a Flat spectrum disk based on J, H, and K band. The source is close to a Class I binary system, Oph-emb 14 VLA 1&2. In the CAMPOS survey, we assume the same bolometric temperature and class for all sources in the close multiple systems. a The SED and protostellar system properties are available in NASA/IPAC Infrared Science Archive. eHOPS classified the source as Class I, but given its bolometric temperature, it should be Class 0. b The eHOPS group calculated the spectral index between Spitzer/IRAC 4.5 $\mu$m and Spitzer/IRAC 24 $\mu$m, and follows Greene et al. (1994) for Class classification (Pokhrel et al. 2023). ODISEA C4 081 or eHOPS-oph-22 shows a negative slope between these wavelengths resembling a Class II by definition. However, the source also shows significant emission at longer wavelengths, resulting in a lower bolometric temperature. c In this paper, if the source has no corresponding eHOPS match, we adopted the non-extinction corrected bolometric temperature from Dunham et al. (2015). d eHOPS source labeled off center for IRAS $1-4\,\mu$m, eHOPS $T_{\rm bol}$ could be wrong, assigned values from Dunham et al. (2015). e eHOPS group updated the SED fit for this source recently, the original values published in Hsieh et al. (2024) for this source were outdated. eHOPS ${\rm T}_{\rm bol}$: Bolometric temperature from eHOPS catalog. The bolometric temperature is derived from SED fitting with Herschel Space Telescope longer wavelength data, but IR extinction correction is not applied. ${\rm L}_{\rm bol}$: Bolometric luminosity. Mass: Disk dust mass. R90%: Disk dust $R_{90\%}$ radius. dist: Distance of the source. $F_{1.3mm}$: 1.3 mm flux. Type: Type of the disk substructure. NA represents no disk substructure is detected. C represents a ring with a large central cavity. F represents the disk is centrally filled. Dun+2015 ${\rm T}_{\rm bol}$: Bolometric temperature from Dunham et al. (2015) catalog. The bolometric temperature is derived from SED fitting without Herschel Space Telescope longer wavelength data, but IR extinction correction is applied. $T_{\rm br,P}$: Brightness temperature of the disk calculated from the full Planck equation. $T_{\rm dust}$: Dust temperature of the disk adopting equation 1. Ref: References. References: (1) Sources observed in our CAMPOS survey. See also the CAMPOS data paper, Hsieh et al. (2024), for source identification and radius measurements. (2) Murillo et al. (2013) (3) Hsieh et al. (2020) (4) Maureira et al. (2022) (5) Ohashi et al. (2023) (6) Narayanan et al. (2023) (7) Cieza et al. (2019) (8) Cieza et al. (2021) (9) Segura-Cox et al. (2020) (10) Andrews et al. (2018a) (11) Huang et al. (2018) (12) ALMA Partnership et al. (2015) (13) Wu et al. (2018)
@@
ODESIA ID SSTgbs/SSTc2d Names eHOPS Name Other Name Included? Notes
@@
A.1 Bolometric temperature as a proxy for relative evolutionary age
Tracing protostellar disk age is extremely difficult. For embedded protostars, the luminosity is dominated by accretion, resulting in a large spread in observed luminosity. This means individual protostellar ages cannot be directly determined from traditional isochrome fitting. Instead, the age sequence for protostellar disks is based on the dissipation of the protostellar envelope, as indicated by observables such as bolometric temperature ($T_{\rm bol}$) (Chen et al. 1995; Myers & Ladd 1993).
@@ A.1 Bolometric temperature as a proxy for relative evolutionary age
Note that this commonly used number-counting method does not take into account different star-formation rates in different molecular clouds, assumes a steady state constant star formation rate, and treats the duration of each stage as a single value rather than a distribution (Kristensen & Dunham 2018). Evans et al. (2009) developed the idea that the protostellar lifetimes should be thought of as half-lives, in which they represent the timescales at which half of the objects will have moved through and left that $T_{\rm bol}$ range. Kristensen & Dunham (2018) accounting for non-steady state star formation rates and adopted a sequential nuclear decay model to estimate the duration of each protostellar stage, and found that the half-life of Class 0 and Class I are $\sim 4.7\times 10^{4}$ and $\sim 8.7\times 10^{4}$ years, respectively.

### v4

@@ 6 Conclusion
In summary, our ALMA study reveals that by $T_{\rm bol}\sim 200-400\,K$ protostellar disk substructure detection rates increase sharply to $\sim$60 %. Our statistical analysis confirms that substructures similar to those in Class II disks are already common by the Class I stage. We argue that this emergence is likely only an upper age limit, as our study suggests increasing optical depth toward the earliest stages.
The prevalence of substructures in the early protostellar stage suggests that planetesimals and giant planets form much more efficiently and earlier than predicted by traditional models (e.g., Pollack et al. (1996)). This early planet formation scenario is also supported by a low abundance of water in the inner Solar System (Morbidelli et al., 2016; Hyodo et al., 2019), the existence of several isotopic reservoirs in the Solar system (Kruijer et al., 2017) as well as the distribution of calcium-aluminum-rich inclusions (Desch et al., 2018; Brasser & Mojzsis, 2020). Additionally, older protoplanetary disks (around Class II sources) might not have enough dust mass to form planetary systems (Manara et al., 2018). Our study has pushed the onset of planet formation to the much earlier protostellar phase, in $\sim$0.2–0.4 Myr disks. During this time period, the central protostar still has a high accretion rate and is building up its mass.
In addition, by classifying disks with substructures into those with (C-type) and without (F-type) central cavities, we find that both populations coexist across evolutionary stages. If protostellar disk substructures are evolutionarily linked, then the data suggest that disk substructures evolve very rapidly and thus can be present at both Class I and Class II stages, and/or that they can be triggered at different times. Alternatively, the observed diversity could be explained by multiple distinct populations of disk substructures, shaped by different formation mechanisms.
@@
All data used in this study are public and can be downloaded from the ALMA Science Archive. LABEL:table:CAMPOS_Tbol, LABEL:table:substructure_data, and LABEL:table:ODISEA2019_Tbol are also available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
C.H.H. and H. A are supported in part by NSF grants AST-1714710. C.H.H. is supported by the NASA Hubble Fellowship Program under award HF2-51556. D.S.C. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2102405. Special thanks to the e-HOPS group, Dr. Riwaj Pokhrel, and Dr. Thomas Megeath for providing the unpublished e-HOPs catalog for protostar classification. C.H.H. thanks Dr. Sean Andrews for the suggestion of considering the effective resolution in constructing a homogeneous sample for this statistical study. We also want to thank the referee for the insightful comments and discussion. This paper makes use of the following ALMA data: ADS/JAO. ALMA #2016.1.00484.L, #2018.1.00028.S, #2019.1.00261.L, #2019.1.01792.S. ALMA is a partnership of the ESO (representing its member states),NSF(USA),and NINS (Japan), together with the NRC (Canada), NSC and ASIAA(Taiwan), and KASI(Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by the ESO, AUI/NRAO, and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This paper makes use of the following software: Astropy: Greenfield et al. (2013); Astropy Collaboration et al. (2018, 2022), CASA: McMullin et al. (2007), SciPy: Virtanen et al. (2020).
@@ Appendix A Supplementary materials
In Table LABEL:table:substructure_data, we present all the disks in this study compiled from the CAMPOS survey (Hsieh et al. 2024), eDisk survey (Ohashi et al. 2023), DSHARP survey (Andrews et al. 2018a), and ODISEA survey (Cieza et al. 2021, 2019). The protostellar systems in this study have bolometric temperatures less than 1900 K, and inclination angles less than $75^{\circ}$. The Table consists of a homogeneous survey of 5 molecular clouds, Corona Australis, Chamaeleon I & II, Ophiuchus & Ophiuchus North, with additional sources in Taurus, Lupus I and BHR 71 from the literature.
In Table LABEL:table:CAMPOS_Tbol, we cross-match all the Corona Australis, Chamaeleon I & II, Ophiuchus North, and Ophiuchus sources in our CAMPOS survey with the Young Stellar Objects catalog from the Spitzer Space Telescope “cores to disks” (c2d) and “Gould Belt” (GB) Legacy surveys (Dunham et al. 2015), as well as the Extension of HOPS Out to 500 ParSecs (eHOPS) catalog to obtain the bolometric temperature ($T_{\rm bol}$), which serves as a proxy for relative evolutionary age for embedded protostellar systems. The eHOPS catalog contains 1-850 $\mu$m SEDs assembled from 2MASS, Spitzer, Herschel, WISE, and JCMT/SCUBA-2 data. It represents the latest and most reliable SED fitting to date. The first eHOPS paper, covering the Serpens and Aquila molecular clouds, was published by Pokhrel et al. (2023). For all other clouds, the SED and the protostellar system properties are available in NASA/IPAC Infrared Science Archive.
@@
CFHTWIR-Oph 43 Flat 440a 570 0.12 0.38 4.4 144 2.0 NA 130 x 100 16.4 23.9 25.3 1
Oph-emb 23 Flat 440a 570 0.12 8.68 50.8 144 46.2 NA 130 x 100 16.4 7.4 25.3 1
Oph-emb 26a Flat 500a 620 0.97 1.9 27.8 144 10.1 NA 130 x 100 16.4 6.3 42.7 1
Oph-emb 26b Flat 500a 620 0.97 1.89 27.7 144 10.1 NA 130 x 100 16.4 6.3 42.7 1
F-MM7 Flat 570a 690 0.72 1.17 27.0 144 6.2 NA 130 x 100 16.4 5.2 39.6 1
Oph-emb 27 Flat 570a 690 0.72 13.0 24.9 144 69.0 NA 130 x 100 16.4 25.2 39.6 1
Oph-emb 28 Flat 610a 720 1.1 2.13 10.3 144 11.3 NA 130 x 100 16.4 24.3 44.0 1
@@
ISO-ChaI 101 Flat 650a 820 0.09 12.06 73.8 179 41.5 C 99 x 65 14.4 6.0 23.6 1
V* DK Cha II 660a 1500 30.0 210.59 56.7 181 708.7 NA 146 x 82 19.8 68.2 100.6 1
V* HO Cha a Flat 710a 900 0.97 4.35 20.4 179 15.0 NA 120 x 90 18.6 14.8 42.7 1
V* HO Cha b Flat 710a 900 0.97 4.43 22.0 179 15.3 NA 120 x 90 18.6 13.5 42.7 1
ISO-ChaI 204 II 720a 1100 0.18 0.46 11.9 179 1.6 NA 120 x 90 18.6 7.2 28.0 1
ISO-ChaI 207 II 1100a 1500 0.09 8.75 90.7 179 30.1 NA 100 x 65 14.4 4.4 23.6 1
ISO-ChaI 237 II 1100a 1600 0.37 8.9 25.8 179 30.6 NA 120 x 90 18.6 17.7 33.5 1
Ass Cha T 1-15 II 1400a 1700 0.45 12.7 46.5 179 43.7 NA 100 x 65 14.4 10.1 35.2 1
@@
111111 CFHTWIR-Oph 79 is also known as GY 263 or SKS 3-48 (Wilking et al. 2015; Narayanan et al. 2023). The classification of this source has been debated. Narayanan et al. (2023) and Allen et al. (2002) classified this source as a Class II disk, while Wilking et al. (2015) classified it as a Flat spectrum disk based on J, H, and K band. The source is close to a Class I binary system, Oph-emb 14 VLA 1&2. In the CAMPOS survey, we assume the same bolometric temperature and class for all sources in the close multiple systems. a In this paper, if the source has no corresponding eHOPS match, we adopt the non-extinction corrected bolometric temperature from Dunham et al. (2015). b The SED and protostellar system properties are available in NASA/IPAC Infrared Science Archive. eHOPS classified the source as Class I, but given its bolometric temperature, it should be Class 0. c eHOPS photometry for this source is not centered at the coordinates of the source for the IRAS $1\,\mu$m image. This could result in the incorrect value of $T_{\rm bol}$. We assigned values from Dunham et al. (2015). d The eHOPS group calculated the spectral index between Spitzer/IRAC 4.5 $\mu$m and Spitzer/IRAC 24 $\mu$m, and follows Greene et al. (1994) for Class classification (Pokhrel et al. 2023). ODISEA C4 081 or eHOPS-oph-22 shows a negative slope between these wavelengths, resembling a Class II by definition. However, the source also shows significant emission at longer wavelengths, resulting in a lower bolometric temperature. e eHOPS photometry for this source is not centered at the coordinates of the source for the IRAS $1-4\,\mu$m image. This could result in the incorrect value of $T_{\rm bol}$. We assigned values from Dunham et al. (2015). f eHOPS group updated the SED fit for this source recently, the original values published in Hsieh et al. (2024) for this source were outdated. eHOPS ${\rm T}_{\rm bol}$: Bolometric temperature from eHOPS catalog. The bolometric temperature is derived from SED fitting with Herschel Space Telescope longer wavelength data, but the IR extinction correction is not applied. ${\rm L}_{\rm bol}$: Bolometric luminosity. Mass: Disk dust mass. R90%: Disk dust $R_{90\%}$ radius. dist: Source distance. $F_{1.3mm}$: 1.3 mm flux. Type: Type of the disk substructure. NA represents that no disk substructure is detected. C represents a ring with a large central cavity. F indicates that the disk is centrally filled. Dun+2015 ${\rm T}_{\rm bol}$: Bolometric temperature from Dunham et al. (2015) catalog. The bolometric temperature is derived from SED fitting without Herschel Space Telescope data, but the IR extinction correction is applied. $T_{\rm br,P}$: Brightness temperature of the disk calculated from the full Planck equation. $T_{\rm dust}$: Dust temperature of the disk adopting equation 1. Ref: References. References: (1) Sources observed in our CAMPOS survey. See also the CAMPOS data paper, Hsieh et al. (2024), for source identification and radius measurements. (2) Murillo et al. (2013) (3) Hsieh et al. (2020) (4) Maureira et al. (2022) (5) Ohashi et al. (2023) (6) Narayanan et al. (2023) (7) Cieza et al. (2019) (8) Cieza et al. (2021) (9) Segura-Cox et al. (2020) (10) Andrews et al. (2018a) (11) Huang et al. (2018) (12) ALMA Partnership et al. (2015) (13) Wu et al. (2018)
@@
ODISEA ID SSTgbs/SSTc2d Names eHOPS Name Other Name Included? Note
@@
Appendix B Bolometric temperature as a proxy for relative evolutionary age
Tracing protostellar disk age is extremely difficult. For embedded protostars, the luminosity is dominated by accretion, resulting in a large spread in observed luminosity. This means individual protostellar ages cannot be directly determined from traditional isochrone fitting. Instead, the age sequence for protostellar disks is based on the dissipation of the protostellar envelope, as indicated by observables such as bolometric temperature ($T_{\rm bol}$) (Chen et al. 1995; Myers & Ladd 1993).
@@ Appendix B Bolometric temperature as a proxy for relative evolutionary age
Note that this commonly used number-counting method does not take into account different star-formation rates in different molecular clouds, assumes a steady state constant star formation rate, and treats the duration of each stage as a single value rather than a distribution (Kristensen & Dunham 2018). Evans et al. (2009) developed the idea that the protostellar lifetimes should be thought of as half-lives, in which they represent the timescales at which half of the objects will have moved through and left that $T_{\rm bol}$ range. Kristensen & Dunham (2018) accounted for a non-steady state star formation rate and adopted a sequential nuclear decay model to estimate the duration of each protostellar stage, and found that the half-life of Class 0 and Class I are $\sim 4.7\times 10^{4}$ and $\sim 8.7\times 10^{4}$ years, respectively.